# Condensed-space methods for nonlinear programming on GPUs

### v1

@@ 5.4 Benchmark on COPS instances
$n$ $m$ it init lin total it init lin total it init lin total
bearing_400 162k 2k 17 0.14 3.42 4.10 14 0.85 0.07 1.13 14 0.78 0.76 1.74
camshape_6400 6k 19k 38 0.02 0.18 0.29 35 0.05 0.03 0.19 38 0.05 0.04 0.23
elec_400 1k 0.4k 185 0.54 24.64 33.02 273 0.46 0.97 20.01 128 0.48 0.75 4.16
gasoil_3200 83k 83k 37 0.36 4.81 5.81 21 0.54 0.24 1.40 20 0.59 0.21 1.35
marine_1600 51k 51k 13 0.05 0.41 0.50 33 0.38 0.58 1.29 13 0.37 0.12 0.62
pinene_3200 160k 160k 12 0.11 1.32 1.60 21 0.87 0.16 1.52 11 0.90 0.84 2.02
robot_1600 14k 10k 34 0.04 0.33 0.45 35 0.20 0.07 0.76 34 0.21 0.08 0.80
rocket_12800 51k 38k 23 0.12 1.73 2.16 37 0.74 0.06 2.49 24 0.25 1.70 3.12
steering_12800 64k 51k 19 0.25 1.49 1.93 18 0.44 0.06 1.64 18 0.46 0.07 1.83
bearing_800 643k 3k 13 0.94 14.59 16.86 14 3.31 0.18 4.10 12 3.32 1.98 5.86
camshape_12800 13k 38k 34 0.02 0.34 0.54 33 0.05 0.02 0.16 34 0.06 0.03 0.19
elec_800 2k 0.8k 354 2.36 337.41 409.57 298 2.11 2.58 24.38 184 1.81 2.40 16.33
gasoil_12800 333k 333k 20 1.78 11.15 13.65 18 2.11 0.98 5.50 22 2.99 1.21 6.47
marine_12800 410k 410k 11 0.36 3.51 4.46 146 2.80 25.04 39.24 11 2.89 0.63 4.03
pinene_12800 640k 640k 10 0.48 7.15 8.45 21 4.50 0.99 7.44 11 4.65 3.54 9.25
robot_12800 115k 77k 35 0.54 4.63 5.91 33 1.13 0.30 4.29 35 1.15 0.27 4.58
rocket_51200 205k 154k 31 1.21 6.24 9.51 37 0.83 0.17 8.49 30 0.87 2.67 10.11
steering_51200 256k 205k 27 1.40 9.74 13.00 15 1.82 0.19 5.41 28 1.88 0.56 11.31
Table 4: COPS benchmark , solved with a tolerance tol=1e-6
@@ 6 Conclusion
This article moves one step further in the solution of generic nonlinear programs on GPU architectures. We have compared two approaches to solve the KKT systems arising at each interior-point iteration, both based on a condensation procedure. Despite the formation of an ill-conditioned matrix, our theoretical analysis shows that the loss of accuracy is benign in floating-point arithmetic, thanks to the specific properties of the interior-point method. Our numerical results show that both methods are competitive to solve large-scale nonlinear programs. Compared to the state-of-the-art HSL linear solvers, we achieve a 10x speed-up on large-scale OPF instances and quasi-dense instances (elec). While the results are more varied across the instances of the COPS benchmark, our performance consistently remains competitive with HSL.
Looking ahead, our future plans involve enhancing the robustness of the two condensed KKT methods, particularly focusing on stabilizing convergence for small tolerances (below $10^{-8}$). It is worth noting that the sparse Cholesky solver can be further customized to meet the specific requirements of the interior-point method wright1999modified . Enhancing the two methods on the GPU would enable the resolution of large-scale problems that are currently intractable on classical CPU architectures such as multiperiod and security-constrained OPF problems.
7 Acknowledgements

### v2

@@ 6.3 Benchmark on CUTEst instances
n m nnzj nnzh it init AD lin total it init AD lin total it init AD lin total
GPU-accelerated solvers are significantly slower
A5NSDSIL 60012 20004 115012 100020 265 0.03 1.22 6.97 11.33 119 26.16 1.35 129.95 164.00 - - - - -
A0NNDNIL 60012 20004 115003 100020 97 0.02 0.43 2.22 3.15 470 17.39 4.10 365.43 395.94 97 17.18 0.93 73.82 93.57
A0NSDSIL 60012 20004 115005 100020 100 0.02 0.43 2.23 3.15 212 18.41 1.88 168.46 192.80 100 18.50 1.20 79.08 100.67
A0NNSNSL 60012 20004 111099 100020 31 0.04 0.14 1.36 1.62 36 19.00 0.34 38.76 58.78 38 19.24 0.48 80.65 101.15
CMPC10 1530 2351 16538 357 70 0.00 0.02 0.24 0.27 82 0.10 0.29 1.93 4.90 70 0.04 0.06 0.78 1.07
CMPC12 1530 2351 16538 357 78 0.00 0.02 0.27 0.31 99 0.21 0.79 4.81 13.11 77 0.04 0.07 0.82 1.14
MPC9 1530 2351 16538 357 70 0.00 0.02 0.25 0.29 79 0.21 0.63 3.25 9.78 69 0.04 0.07 0.75 1.05
LUKVLE8 10000 9998 29994 30000 12 0.01 0.05 0.04 0.10 144 0.12 1.36 2.48 25.83 12 0.04 0.05 13.45 13.57
READING7 1002 500 126251 1503 125 0.01 0.14 1.28 1.49 330 0.65 1.45 4.75 18.18 120 0.55 0.35 1.25 3.09
BRAINPC0 6907 6900 41399 17335 120 0.01 0.27 0.78 1.11 43 0.22 0.47 2.83 7.65 - - - - -
ROCKET 2407 2002 10405 8019 58 0.00 0.05 0.10 0.16 38 0.21 0.35 1.71 5.11 - - - - -
HYDROELL 1009 1008 2016 2017 235 0.00 0.03 0.08 0.15 235 0.05 0.41 0.70 4.65 235 0.02 0.16 0.19 0.95
DALE 16514 405 33028 16514 14 0.01 0.01 0.07 0.11 16 2.12 0.05 30.59 33.03 14 1.66 0.03 16.60 18.36
CBS 11163 244 22326 11163 20 0.00 0.01 0.07 0.10 22 0.75 0.03 7.60 8.49 35 0.74 0.07 8.92 9.91
BLOCKQP2 10010 5001 70010 15010 19 0.01 0.02 0.09 0.13 37 36.27 0.08 7.37 44.62 19 36.30 0.05 2.75 39.50
ROSEPETAL2 10001 20001 50001 10000 29 0.05 0.04 0.15 0.29 47 36.11 0.17 7.41 44.86 29 36.15 0.17 3.24 40.16
GPU-accelerated solvers are significantly faster
STCQP1 8193 4095 28665 57333 13 0.01 0.06 13.87 13.96 15 0.06 0.08 0.06 0.26 161 0.07 1.10 1.18 2.94
CVXQP3 10000 7500 22497 39984 17 0.12 0.03 36.33 36.50 31 0.13 0.07 0.41 0.69 17 0.13 0.07 118.33 118.60
CATENA 3003 1000 6000 6003 32 0.00 0.02 12.72 12.74 36 0.06 0.10 0.33 1.13 - - - - -
CLEUVEN3 1200 2973 199600 71583 134 0.08 0.80 6.81 7.87 104 0.53 0.80 0.29 2.16 99 0.54 0.84 1.07 3.00
NCVXQP7 10000 7500 22497 39984 121 0.12 0.21 133.73 134.46 302 0.13 1.12 7.91 10.55 129 0.13 0.48 178.30 179.46
BDRY2 251001 250498 1247999 63500 45 2.53 1.06 158.10 163.68 157 1.46 4.99 28.96 36.30 50 1.75 1.57 10.99 14.77
SCW2 2003 1001 2002 2007006 251 0.41 25.96 66.30 95.12 97 0.75 10.64 1.55 13.48 259 0.74 28.03 4.71 35.14
CLEUVEN6 1200 3091 216608 71583 206 0.08 1.26 9.45 11.05 88 0.55 0.74 0.27 2.04 - - - - -
AUG3D 27543 8000 50286 22743 2 0.16 0.01 1.45 1.61 3 0.22 0.01 0.04 0.29 1 0.22 0.01 0.03 0.26
ROBOTARM 4412 3202 15605 8422 451 0.00 0.28 1.38 1.81 36 0.06 0.04 0.18 0.46 - - - - -
EIGENB2 2550 1275 125000 66300 132 0.08 0.82 211.36 212.42 110 1.23 1.01 1.77 4.68 128 1.23 1.21 2.14 5.25
CBRATU3D 3456 2000 16000 3000 3 0.03 0.00 0.29 0.32 5 0.06 0.01 0.03 0.13 3 0.07 0.01 0.03 0.12
EIGENC2 2652 1326 132651 70278 16 0.08 0.10 26.37 26.57 17 1.31 0.16 0.58 2.15 15 1.32 0.16 0.56 2.12
STNQP1 8193 4095 28665 57333 17 0.01 0.07 19.59 19.69 21 0.06 0.12 0.13 0.38 60 0.06 0.40 0.48 1.24
EIGENCCO 2652 1326 132651 3517878 42 0.58 8.85 51.38 61.37 35 1.46 7.56 0.59 9.84 42 1.77 9.09 0.83 11.93
EIGENAU 2550 2550 313750 66300 2 0.05 0.03 10.95 11.05 3 1.49 0.05 0.02 1.60 1 1.50 0.03 0.01 1.57
LUKVLE12 9997 7497 19992 22492 86 0.00 0.18 208.54 208.77 30 0.05 0.09 0.09 0.41 24 0.04 0.07 0.08 0.30
MSQRTB 1024 1024 64512 33264 5 0.01 0.01 1.62 1.65 6 0.23 0.02 0.03 0.30 5 0.23 0.02 0.02 0.29
DRCAVTY1 4489 3969 51597 86879 8 0.01 0.07 1.19 1.28 8 0.08 0.09 0.05 0.25 7 0.08 0.09 0.05 0.24
CLEUVEN5 1200 2973 199600 71583 134 0.07 0.81 6.80 7.85 104 0.53 0.82 0.29 2.18 99 0.53 0.77 1.11 3.04
MARINE 11215 11192 46768 9423 85 0.01 0.11 8.60 8.75 25 0.09 0.05 0.14 0.39 31 0.10 0.07 0.25 0.54
Table 4: CUTEst benchmark, solved with a tolerance tol=1e-6. The derivatives are evaluated with CUTEst. The column init displays the time spent initializing the IPM solver, including the initial symbolic factorization in the linear solver. The columns AD, lin and total display the time spent in the evaluation of the derivatives using AD, the time spent in the linear solver and the total time spent in the IPM algorithm.
@@ 7 Conclusion
This paper investigates the theoretical and numerical properties of two condensed-space interior-point methods for solving large-scale nonlinear programming problems on GPUs. Despite the emergence of ill-conditioned matrices, our theoretical analysis shows that the resulting loss of accuracy is benign in floating-point arithmetic, owing to the structural properties of the interior-point method. Numerically, both methods prove competitive for solving large-scale nonlinear problems. The fully GPU-resident software stack we developed, which includes MadNLP.jl, cuDSS, and ExaModels.jl, demonstrates its effectiveness on large-scale optimal power flow (OPF) problems from the pglib-opf library. Notably, this GPU-based implementation achieves up to a 10× speedup on large-scale OPF problems. While performance is more variable across the CUTEst benchmarks, our solver consistently remains competitive with its CPU-based counterpart. When focused on the instances with non-trivial size, the GPU implementation achieves considerable speedups, at the price of reduced robustness in the linear solver. These results highlight both the potential (high performance and scalability) and the limitations (robustness and convergence) of GPU-based interior-point methods for nonlinear programming. Looking ahead, we plan to improve the robustness of the two condensed KKT methods, with a particular emphasis on achieving stable convergence at tighter tolerances (below $10^{-8}$). Moreover, the sparse Cholesky solver can be further customized to align with the specific structure and needs of interior-point algorithms wright1999modified .
Fundings